OpenStudy (pig_cat223):

please help me!!! solve this system of linear equations. separate the x- and y-values with a comma. -16x=50-14y -7x=77-14y

OpenStudy (theprofessor35):

A system can be solved by substitution or elimination.

OpenStudy (theprofessor35):

Lets do substitution method.

OpenStudy (theprofessor35):

Okay... so... -16x=50-14y -7x=77-14y -16x = 50 - 14y x = 50/-16 - 14y/-16

OpenStudy (theprofessor35):

Simplify that to get....

OpenStudy (theprofessor35):

x = 50/-16 - 14y/-16 x = - 3.125 + 14/16y

OpenStudy (theprofessor35):

Now plug that into the other equation and solve for y.

OpenStudy (theprofessor35):

-7x=77-14y -7(- 3.125 + 14/16y ) = 77 - 14y 21.875 - 6.125y = 77 - 14y

OpenStudy (theprofessor35):

21.875 - 6.125y = 77 - 14y +6.125y +6.125y

OpenStudy (theprofessor35):

21.875 = 77 - 7.875y -77 -77

OpenStudy (andrewkaiser333):

so -55.125=-7.875y

OpenStudy (andrewkaiser333):

right?

OpenStudy (theprofessor35):

Solve for y now. and then plug the y value into the OTHER equation to solve for x.

OpenStudy (theprofessor35):

and then you shall have the system's solution! :)

OpenStudy (pig_cat223):

thanks guys... :D

OpenStudy (theprofessor35):

Welcome, please medal if you would like. :)